OpenStudy (anonymous):

Solve. One point for showing the steps to solve each equation and one point for the correct solution. 4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

4x - 24 = -2x + 6 4x + 2x= 6+24 6x = 30 x=5